Stephen plans to purchase a car 6 years from now. The car will cost $38,586 at that time. Assume that Stephen can earn 9.40 percent (compounded monthly) on his money. How much should he set aside today for the purchase?

Round the answer to two decimal places.
Answer #1

Let amount to be set aside be $ x

rate per month for compounding = r = 9.4%/12 = 0.783333%

n = number of months = 12*6 = 72

therefore $38,586 = x * (1+0.0078333)^72

x = $38,586 / 1.75383 = $22,000.99

So stephen should set aside $22,000.99 today for purchase